OpenStudy (anonymous):

HELP, PLEASE. 3(x-2)≤1/3 solve inequality

OpenStudy (anonymous):

this is an "and" case we can set up a 3 part inequality

OpenStudy (anonymous):

I am confused what you did there. Did you divide 1/3 to both sides?

OpenStudy (anonymous):

correct divide 3 by both sides

OpenStudy (anonymous):

what did you get

OpenStudy (anonymous):

3(x-2)/3 <=1/3 /3 Would I distribute first and then divide or just divide the whole thing?

OpenStudy (anonymous):

You distribute first

OpenStudy (anonymous):

I was thinkingg about absolute value sorry

OpenStudy (anonymous):

3x-6<=1/3

OpenStudy (anonymous):

add 6

OpenStudy (anonymous):

To both sides?

OpenStudy (anonymous):

or divide by 3

OpenStudy (anonymous):

3x<= 6 1/3

OpenStudy (anonymous):

yes what you do to one side you do to the other

OpenStudy (anonymous):

Then I would divide by three?

OpenStudy (anonymous):

6+1/3 18/3 +1/3 19/3

OpenStudy (anonymous):

3x<== 19/3 x<== 19/9

OpenStudy (anonymous):

i got x 19/9

OpenStudy (anonymous):

Did you turn it into an improper fraction? I am unsure how you got 18/3 + 1/3

OpenStudy (anonymous):

* x==>19/9

OpenStudy (anonymous):

yes i turned 6 into an improper fraction

OpenStudy (anonymous):

6= 18/3

OpenStudy (anonymous):

Okay so 19/3 <= X?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

x has to bes less than or equal to 19/9

OpenStudy (anonymous):

you forgot to divide it by 3

OpenStudy (anonymous):

Right, I was about to mention.

OpenStudy (anonymous):

So we have 3x => 6 1/3 divided by 3?

OpenStudy (anonymous):

so we get x==> 2 1/9

OpenStudy (anonymous):

2 1/9 = 19/9

OpenStudy (anonymous):

Right, so X<= 19/9 or 2 1/9

OpenStudy (anonymous):

no x==> 19/9 we did not divide by a negative to change the sign

OpenStudy (anonymous):

So X is greater than or equal to 19/9?

OpenStudy (anonymous):

\[x \le 19/9\]

OpenStudy (anonymous):

less than

OpenStudy (anonymous):

or equal to

OpenStudy (anonymous):

Okay, thanks a lot for walking me through it.

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

Could you help me verify my solution to using some of the numbers from the solution set?

OpenStudy (anonymous):

so it has to be less than 2 1/9 lets try 0 since its the easiest

OpenStudy (anonymous):

f(0) = 3(0-2)≤1/3 3(-2) ≤1/3 -6≤1/3 its checks out

OpenStudy (anonymous):

Awesome, thanks!